OpenStudy (anonymous):

Anyone know a way to factor 5x^2 + 34x - 7

OpenStudy (anonymous):

(5)(-7)=-35 then since -35*1=-35 and -35+1=34 we can rewrite 34x\[=5x^2-35x+x-7\]now we factor by grouping\[=5x(x-7)+(x-7)\]\[=(x-7)(5x+1)\]this is the factored form

OpenStudy (anonymous):

Ah okay :D tyyy

OpenStudy (anonymous):

np :})

OpenStudy (anonymous):

you're awesome